# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v1

@@ III-A4 MAHNOB-HCI
III-B Data Preprocessing
III-C Experimental Task
The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications. The cross-session task necessitates datasets with multiple sessions, which many datasets lack. Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance. Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects.
Figure 2: The framework and data pipeline of LibEER
III-D Data Splitting
III-E Evaluation Methods
@@
Therefore, in our benchmark, our evaluation method reports the performance on the test set of the model that achieved the highest F1 score on the validation set across all epochs. We use accuracy and F1 score, the most commonly used metrics in existing research, and report both the mean and standard deviation. The methods for calculating accuracy and F1 score are presented in Eq. 5 and Eq. 6, respectively.
@@ IV LibEER Toolkit Framework
In this section, we will present the framework of the algorithm library to facilitate researchers’ utilization of these models in the domain of EER. Specifically, we will detail the coding procedures associated with the data loader, data split, and model training and evaluation within the LibEER. To facilitate the utilization of the LibEER library in the field of EER, users can easily install via pip, which stremlines the installation process and ensures that all necessary dependencies are managed automatically. For further details and access to the source code, please visit the official repository at https://github.com/ButterSen/LibEER.
LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch. It comprises three primary modules: the data loader, data splitting, and model training and evaluation. The data loader module standardizes the data formats across various datasets and provides a range of preprocessing techniques. The data splitting module delivers a cohesive solution applicable to different experimental tasks and partitioning methods. The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation. Figure 2 illustrates the framework of LibEER. Below, we provide detailed explanations and usage examples for each module.
@@ IV-1 Data loader
Additionally, LibEER provides the preprocess() function, which allows for the preprocessing of data based on user-defined settings. The preprocess() function operates on a trial basis and encompasses three primary steps: noise removal, feature extraction, and sample segmentation. When utilizing raw data directly from the dataset, only sample segmentation is executed. Noise removal aims to eliminate electromyography (EMG) noise, electrooculography (EOG) artifacts, and other forms of interference from the EEG signals. The feature extraction process analyzes the data to derive relevant EEG features, such as DE and PSD features, based on the specified feature types and frequency bands provided by the user. Sample segmentation organizes the data into user-defined sample lengths. Ultimately, the processed data is integrated into the format of (session, subject, trial, sample). This functionality is illustrated in the usage examples of LibEER for data loading. A data loader usage example is presented in Figure 3.
@@ IV-2 Data split
To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels. The splitting methodologies available in LibEER include fixed ratio splitting, which partitions the data into training and testing sets according to a predetermined ratio; early stopping splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution; and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once. The following example (see Figure 4) illustrates how to employ LibEER for data splitting.
@@ IV-3 Model training and evaluation
The ultimate outcomes of the model will be reported according to the performance metrics specified by the user. The following example in Figure 5 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
V Experimental Validation
@@
TABLE V: The Mean Accuracies (and Standard Deviations) using proposed benchmark for Subject-Dependent EEG Emotion Recognition Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
@@
TABLE VI: The Mean Accuracies using proposed benchmark for Cross-Subject EEG Emotion Recognition Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
@@ V-A Model Reproduction
To ensure the reliability of our benchmark, we conducted a meticulous replication of the representative EER methods using the original configurations and compared them with the results reported in the original paper. Table IV presents both our reimplementation outcomes and those reported in the original study with the gap. Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models.
@@
(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality. Notably, the reproduction results for RGNN, DBN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper. Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study. The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers. We determine these parameters and settings through tuning and experience, which could introduce some biases into the results. Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
@@ V-B Model Comparison
As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field. To this end, we conduct a fair comparison of all methods based on the unified and reasonable benchmark proposed in Section III. The results for the two experimental tasks are reported in Tables V and VI, respectively. We select SVM [28], one of the most commonly used machine learning algorithms, as the baseline model. Its results are presented first. In the two tables, the top two methods in each scenario are highlighted using bold and underlined formatting. From these results, we make the following key observations.
@@
To provide a clearer illustration of the performance of each method under the benchmark, we ranked each method’s performance in each scenario. The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point. The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Figure 6 (a) and (b).
From Figure 6 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140. Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points. CDCN and RGNN ranked fourth and fifth with scores of 126 and 105, respectively. In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly. In the cross-subject task (see Figure 6 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario. Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario. In contrast, the MS-MDA model, specifically designed for cross-subject scenarios, shows significant variability across different datasets, resulting in a moderate ranking in the overall score.
@@
To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Figure 7. From the results presented in Figure 7, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI. We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective. Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not. However, as the number of labels increases, all models demonstrate similar performance on the four-class tasks of SEED-IV, DEAP-VA, and HCI-VA, with relatively poorer performance observed in DEAP-VA.
@@
V-C Discussion of Ablation Study
@@
V-C1 Sample length
In the ablation study on sample length, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Figure 8. As illustrated in Figure 8, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases. We attribute this trend to the sparsity of data in the subject-dependent context; shorter sample lengths yield a greater total number of samples, which enhances the learning capacity of deep models.
@@ V-C1 Sample length
V-C2 Feature types
In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs. The results are presented in Figure 9. As illustrated in Figure 9, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements. In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy. This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
V-C3 Training Methods
@@
VI Limitation
While LibEER is able to standardize and accelerate research in EER, we are aware of some potential limitations described as follows:
Incomplete Experimental Settings. LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods. However, LibEER does not cover all experimental configurations used in the EER field. Some studies use other datasets like DREAMER[46], Amigos[47], and MPED[48]. Others develop unique preprocessing methods, including customized filtering, artifact removal[8, 51, 52], denoising[49, 50], normalization techniques[50, 8], and extracting other types of features [12, 44, 14]. Additionally, some tasks, such as cross-session[53, 54, 55], subject-independent[54, 37], and cross-dataset tasks[45, 50], are designed to assess model performance under different conditions, such as data from different sessions or populations. Furthermore, alternative evaluation strategies, including early stopping[56, 57, 58], are employed in some studies for more robust model training and assessment. Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.
Inadequate Model Reproduction. In replicating representative models in the EER field, we faced several challenges despite closely following the papers. Some methods lacked detailed descriptions of model or training parameters[5, 7]. In other cases [12, 13, 11], the available code did not fully align with the reported paper details. We also found that in DGCNN[12] minor adjustments during replication improved performance, suggesting the paper’s content may not have been updated. When the above issues arose, we used our experience to replicate the models as accurately as possible, adhering to the papers. The replication results (Table IV) show discrepancies compared to the original reports. In the future, we aim to collaborate with original authors and the broader EER community to improve and expand this algorithm library.

### v2

@@ III-D4 MAHNOB-HCI
III-E Evaluation Methods and Metrics
@@
Therefore, in our benchmark, our evaluation method reports the performance on the test set of the model that achieved the highest F1 score on the validation set across all epochs. We report both the mean and standard deviation of two metrics, accuracy and F1 score. The former is a key metric for classification tasks, while the latter provides a more reasonable assessment when sample labels are imbalanced. The methods for calculating accuracy and F1 score are presented in Eq. 6 and Eq. 7, respectively.
@@ IV LibEER Toolkit Framework
In this section, we will present the framework of the algorithm library to facilitate researchers’ utilization of these models in the domain of EER. Specifically, we will detail the coding procedures associated with the data loader, data split, and model training and evaluation within the LibEER. To facilitate the utilization of the LibEER library in the field of EER, users can easily install via pip, which stremlines the installation process and ensures that all necessary dependencies are managed automatically. For further details and access to the source code, please visit the official repository at https://github.com/XJTU-EEG/LibEER.
LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch. It comprises three primary modules: the data loader, data splitting, and model training and evaluation. The data loader module standardizes the data formats across various datasets and provides a range of preprocessing techniques. The data splitting module delivers a cohesive solution applicable to different experimental tasks and partitioning methods. The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation. Fig. 3 illustrates the framework of LibEER. Below, we provide detailed explanations and usage examples for each module.
@@ IV-1 Data loader
Additionally, LibEER provides the preprocess() function, which allows for the preprocessing of data based on user-defined settings. The preprocess() function operates on a trial basis and encompasses three primary steps: noise removal, feature extraction, and sample segmentation. When utilizing raw data directly from the dataset, only sample segmentation is executed. Noise removal aims to eliminate electromyography (EMG) noise, electrooculography (EOG) artifacts, and other forms of interference from the EEG signals. The feature extraction process analyzes the data to derive relevant EEG features, such as DE and PSD features, based on the specified feature types and frequency bands provided by the user. Sample segmentation organizes the data into user-defined sample lengths. Ultimately, the processed data is integrated into the format of (session, subject, trial, sample). This functionality is illustrated in the usage examples of LibEER for data loading. A data loader usage example is presented in Fig. 4.
@@ IV-2 Data split
To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels. The splitting methods available in LibEER include train-test splitting, which partitions the data into training and testing sets according to a predetermined ratio;train-val-test splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution;and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once. The following example (see Fig. 5) illustrates how to employ LibEER for data splitting.
@@
BiDANN SEED B, R, DE, 9s dependent 3 : 2 ACC 92.38$\pm$7.04 89.06$\pm$9.42 3.32$\downarrow$
R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 84.11$\pm$8.47 9.27$\downarrow$
@@ IV-3 Model training and evaluation
The ultimate outcomes of the model will be reported according to the performance metrics specified by the user. The following example in Fig. 6 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
V Experimental Results
@@
TABLE V: The Mean Accuracies and F1 scores (and Standard Deviations) using the proposed benchmark for Subject-Dependent EER Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
@@
TABLE VI: The Mean Accuracies and F1 scores using the proposed benchmark for Cross-Subject EER Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
@@ V-A Model Reproduction
To ensure the reliability of our benchmark, we conducted a meticulous replication of the representative EER methods using the original configurations and compared them with the results reported in the original paper. Table IV presents both our reimplementation outcomes and those reported in the original study with the gap. Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models. All reproduced models have been integrated into our LibEER algorithm library, allowing for direct usage.
@@
(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality. Notably, the reproduction results for RGNN, DBN, R2G-STNN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper. Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study. The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers. We determine these parameters and settings through tuning and experience, which could introduce some biases into the results. Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
@@ V-B Model Comparison
As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field. To this end, we conduct a fair comparison of ten selected methods based on the unified and reasonable benchmark proposed in Section III. The results for the two experimental tasks are reported in Tables V and VI, respectively. We select SVM [29], one of the most commonly used machine learning algorithms, as the baseline model. Its results are presented first. In the two tables, the top two methods in each scenario are highlighted using bold and underlined formatting. From these results, we make the following key observations.
@@
To provide a clearer illustration of the performance of each method under the benchmark, we ranked each method’s performance in each scenario. The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point. The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Fig. 7 (a) and (b).
From Fig. 7 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140. Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points. CDCN and RGNN ranked fourth and fifth with scores of 126 and 105, respectively. In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly. In the cross-subject task (see Fig. 7 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario. Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario. In contrast, the MS-MDA model, specifically designed for cross-subject scenarios, shows significant variability across different datasets, resulting in a moderate ranking in the overall score.
@@
To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Fig. 8. From the results presented in Fig. 8, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI. We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective. Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not. However, as the number of labels increases, all models demonstrate similar performance on the four-class tasks of SEED-IV, DEAP-VA, and HCI-VA, with relatively poorer performance observed in DEAP-VA.
@@
V-C Discussion of Experiment Settings
@@
V-C1 Data Splitting Length
Existing studies have chosen to divide trials into samples of 1 to 10 seconds in length [12, 8, 11]. To explore the effects and impact of different sample lengths, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Fig. 9. As illustrated in Fig. 9, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases. We attribute this trend to the sparsity of data in the subject-dependent context; shorter sample lengths yield a greater total number of samples, which enhances the learning capacity of deep models.
@@ V-C1 Data Splitting Length
V-C2 Data preprocessing Feature
In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs. The results are presented in Fig. 10. As illustrated in Fig. 10, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements. In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy. This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
V-C3 Evaluation Methods
@@
VI Future Direction
While LibEER is able to standardize and accelerate research in EER, we will further improve and enhance the following two aspects in our future work.
Model Reproduction. In replicating representative models in the EER field, we faced several challenges despite closely following the papers. Some methods lacked detailed descriptions of model or training parameters[5, 7]. In other cases [12, 13, 11], the available code did not fully align with the reported paper details. We also found that in DGCNN[12] minor adjustments during replication improved performance, suggesting the paper’s content may not have been updated. When the above issues arose, we used our experience to replicate the models as accurately as possible, adhering to the papers. The replication results (Table IV) show discrepancies compared to the original reports. In the future, we aim to collaborate with original authors and the broader EER community to improve and expand this algorithm library.
Experimental Settings. LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods. However, LibEER does not cover all experimental configurations used in the EER field. Some studies use other datasets like DREAMER[47], Amigos[48], and MPED[49]. Others develop unique preprocessing methods, including customized filtering, artifact removal[8, 52, 53], denoising[50, 51], normalization techniques[51, 8], and extracting other types of features [12, 45, 15]. Additionally, some tasks, such as cross-session[54, 55, 56], subject-independent[55, 38], and cross-dataset tasks[46, 51], are designed to assess model performance under different conditions, such as data from different sessions or populations. Furthermore, alternative evaluation strategies, including early stopping[57, 58, 59], are employed in some studies for more robust model training and assessment. Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.